I am working on Systems of Equations.
Jackie bought some stamps for $15.16. Some were 26 cent stamps and soe were 41 cent stamps. She bought 41 stamps in all. How many of each stamp did she buy?
I have defined variables as: a=# of 26 cent stamps and b=# of 41 cent stamps
I have written out the two equations (i believe they are right but they may be the reason I can't figure out the problem): .26a+.41b=15.16
a+b=41
I have multiplied the first one by 100 to eliminate the decimals, but as I go through the problem using elimination ( that is what I was directed to use my my teacher) I end up with decimals and that can't be right as you can't have 2.121212121 stamps...
Hopefully you can help me, it would be much appreciated and thank-you. :)

Answer by scott8148(6628)

26a + 41b = 1516

26(a + b = 41) ___ 26a + 26b = 1066

subtracting (to eliminate a) ___ 15b = 450 ___ b = 30